OpenStudy (nicole143):

Determine whether each sequence is arithmetic. if so, identify the common difference and find the 32nd term. Sequence: 2, 5, 9, 14 I know how to do it when the increase is constant but I don't know how to change it when it goes up by different numbers. @kc_kennylau

OpenStudy (kc_kennylau):

First of all is it arithmetic?

OpenStudy (nicole143):

Yes?

OpenStudy (nicole143):

Or is that only when it goes up by a constant number?

OpenStudy (nicole143):

It's no.. Ha sorry

OpenStudy (kc_kennylau):

You're correct, that's a no :)

OpenStudy (nicole143):

Okay, so different question. When trying to find the missing term and theres more than one, how would the equation change from just have 2 on the bottom?

OpenStudy (kc_kennylau):

Uh... Sorry I don't get what you mean :/

OpenStudy (nicole143):

I have to find the missing term in the sequence 2, ___, ___, ___, -0.4 I did another one like that one^^ but I only had to find one term and the equation was simply 11+27/2 I thought that the dominator corresponded with whatever term you had to find but my idea on that didn't work. How do I find the the answer for this one?

OpenStudy (kc_kennylau):

If you have the first term and the common difference you can generate the whole sequence right?

OpenStudy (kc_kennylau):

I'll denote the first term be "a" and the common difference be "d" from now on

OpenStudy (kc_kennylau):

If you have "a" and "d", the whole sequence is: a, a+d, a+2d, a+3d, ..., a+(n-1)d

OpenStudy (nicole143):

Um, I don't know. My book looks different..

OpenStudy (kc_kennylau):

What does your book say

OpenStudy (the_fizicx99):

The difference is 0.8 :>

OpenStudy (kc_kennylau):

Nope it isn't

OpenStudy (the_fizicx99):

Yes, it is. 2-0.8 = 1.2 1.2 - 0.8 = 0.4 0.4 - 0.8 = -0.4

OpenStudy (nicole143):

Okay, so my the original problem says: Find the missing term(s) of each arithmetic sequence. Sequence: 2, ___, ___, ___, -0.4

OpenStudy (kc_kennylau):

And no it isn't @tHe_FiZiCx99

OpenStudy (kc_kennylau):

So let a be the first term and d be the common difference, n=5 as there are 5 terms

OpenStudy (kc_kennylau):

a=2 a+(5-1)d=-0.4

OpenStudy (kc_kennylau):

Let's solve for d

OpenStudy (nicole143):

Can you please write there equation in the equation maker thing so I can see it? It makes it easier for me..

OpenStudy (kc_kennylau):

What difference does it make? \[a=2\]\[a+(5-1)d=-0.4\]

OpenStudy (nicole143):

I don't know, it's just easier for me to see it clearly. I'm a better visual learner..

OpenStudy (kc_kennylau):

So solve for d?

OpenStudy (nicole143):

Okay, one second.

OpenStudy (the_fizicx99):

It's -0.8 ._. Yes, you have a 3 number space gap so subtract 2 - (-0.4) = 2.4 / 3 = 0.8 a_n = a1 + d(n-1) a_2 = 2 + (-0.8)(2-1) a_2 = 2 + (-0.8)(1) a_2 = 2 - 0.8 a_2 = 1.2 Alternatively you can skip the last digit (4) a_5 = 2 + (-0.8)(4-1) a_5 = 2 + (-0.8)(3) a_5 = 2 + (-2.4) a_5 = -0.4

OpenStudy (nicole143):

d= -0.4 ?

OpenStudy (kc_kennylau):

Nope :/

OpenStudy (the_fizicx99):

That 5 is a typo X_X

OpenStudy (nicole143):

No to me or to him?

OpenStudy (kc_kennylau):

@tHe_FiZiCx99 double check it @Nicole143 to you

OpenStudy (nicole143):

Okay, um. I don't know then.

OpenStudy (kc_kennylau):

How did you solve it? Maybe I can find what's wrong in your steps

OpenStudy (nicole143):

I may have done it wrong from the beginning. 2 + (5+1)d = -0.4 -2 6d = -2.4 / 6 d = -0.4

OpenStudy (the_fizicx99):

Oh, 0.8 skips one dash. Yeah, typo. >.<

OpenStudy (kc_kennylau):

It's 5-1 instead of 5+1

OpenStudy (kc_kennylau):

to Nicole

OpenStudy (nicole143):

Oh, okay. One second.

OpenStudy (nicole143):

D = -0.6

OpenStudy (kc_kennylau):

Yep :)

OpenStudy (kc_kennylau):

So can you find now the second term, the third term and fourth term?

OpenStudy (kc_kennylau):

Hint: The first term is a, the second term is a+d, the third term is a+2d, ...

OpenStudy (nicole143):

Use 3 and 4 instead of 2?

OpenStudy (kc_kennylau):

Nope

OpenStudy (kc_kennylau):

You now have a=2 and d=-0.6

OpenStudy (kc_kennylau):

Find the second term i.e. a+d

OpenStudy (nicole143):

What?

OpenStudy (kc_kennylau):

Find a+d

OpenStudy (nicole143):

-2.6 ??

OpenStudy (kc_kennylau):

What's 2+(-0.6)

OpenStudy (nicole143):

1.4 ...

OpenStudy (kc_kennylau):

Exactly

OpenStudy (nicole143):

So do I then replace 2 for 3 and 4? Then take a+d form those equations for the next two?

OpenStudy (kc_kennylau):

You don't replace 2

OpenStudy (kc_kennylau):

2 is the first term

OpenStudy (nicole143):

What do I replace to get the next 2 term?

OpenStudy (kc_kennylau):

The first term is the first term; The second term is the first term plus the common difference; The third term is the first term plus twice the common difference; The fourth term is the first term plus thrice the common difference; The fifth term is the first term plus four times the common difference.

OpenStudy (kc_kennylau):

\(a_1=a\) \(a_2=a+d\) \(a_3=a+2d\) \(a_4=a+3d\) \(a_5=a+4d\)

OpenStudy (nicole143):

The common difference being -0.6 ?

OpenStudy (kc_kennylau):

Yes

OpenStudy (nicole143):

Okay, so the third term would be .8 and the fourth would be .2 ?

OpenStudy (kc_kennylau):

Exactly

OpenStudy (nicole143):

Great! Thank you!

OpenStudy (kc_kennylau):

Any questions? Asking questions is an effective way of learning

OpenStudy (nicole143):

I don't think so

OpenStudy (the_fizicx99):

>.> I see why I got 0.8, missed adding a 1 to the divisor :/

OpenStudy (kc_kennylau):

@tHe_FiZiCx99 Glad you know where your mistake is.

OpenStudy (kc_kennylau):

@Nicole143 okay :)

OpenStudy (the_fizicx99):

Ahem, was* xD and it's a shortcut to getting the difference lol

OpenStudy (kc_kennylau):

Lol my grammar sucks xp

OpenStudy (kc_kennylau):

And you're right it's a shortcut but still derived from my way

OpenStudy (kc_kennylau):

a=2 and a+(5-1)d=-0.4

OpenStudy (kc_kennylau):

(2)-(1): (5-1)d=-0.4-2

OpenStudy (kc_kennylau):

d=(-0.4-2)/(5-1)

OpenStudy (kc_kennylau):

which is your way

OpenStudy (the_fizicx99):

Mine was actually dividing lol 2 + 0.4= 2.4/4 = 0.6 because it's decreasing you add the negative sign.

OpenStudy (kc_kennylau):

Yep that's just the same :)

OpenStudy (the_fizicx99):

Simpler lol

OpenStudy (nicole143):

Can you help me with another?

OpenStudy (nicole143):

I just tried to do it the same way we did another before but it didn't work..

OpenStudy (kc_kennylau):

I don't have time, so @tHe_FiZiCx99 will help you I suppose?

OpenStudy (nicole143):

Okay, @tHe_FiZiCx99 can you help?

OpenStudy (the_fizicx99):

Um sure

OpenStudy (nicole143):

Okay, the problem is finding the explicit formula for the arithmetic sequence. Sequence: 62, 59, 56, 53 I tried doing the same thing that kc showed me before but it didn't come out with the right answer..

OpenStudy (kc_kennylau):

What was your answer?

OpenStudy (nicole143):

an = 3n + 65

OpenStudy (kc_kennylau):

The common difference is -3 instead of 3 :)

OpenStudy (kc_kennylau):

Common difference = Second-first = Third-second = Fourth-third = ...

OpenStudy (nicole143):

Yes, I did that part

OpenStudy (kc_kennylau):

It should be -3n instead of 3n

OpenStudy (nicole143):

But the book says that the answer should be an=62-3(n-1) How do I get that?

OpenStudy (kc_kennylau):

Oh, I just helped you expand it

OpenStudy (kc_kennylau):

Maybe my way wasn't the standard way :P

OpenStudy (kc_kennylau):

So what's the common difference

OpenStudy (nicole143):

Haha okay, so dod you know how to get to that answer?

OpenStudy (nicole143):

*do